(x+3)^2=64
x^2-16=9

Answer by stanbon(75887)

(x+3)^2=64
---
(x+3)^2-8^2 = 0
---
Factor:
(x+3-8)(x+3+8) = 0
----
(x-5)(x+11) = 0
----
x = 5 or x = -11
====================

x^2-16=9
---
x^2-25 = 0
Factor:
(x-5)(x+5) = 0
---
x = 5 or x = -5
====================
Cheers,
Stan H.
==========